# Machine Translation Models are Zero-Shot Detectors of Translation Direction

### v2

@@ 5.6 Application to Real-World Forensic Case
Finally, we apply our approach to the 86 segment pairs of the plagiarism allegation case. We treat the segments as a single document and classify them with M2M-100 using the document-level approach defined in Section 3.2. We find that according to the model, it is more probable that the English segments are translations of the German segments than vice versa.
@@
Overall, our analysis supports the hypothesis that German is indeed the language of origin in this real-world dataset (forgery hypothesis; Figure 2). Nevertheless, we recommend that additional evidence be considered before drawing a final conclusion, given the error rate of 5–21% that we observed in our earlier experiments on English–German WMT documents.
@@ 6 Conclusion
We proposed a novel approach to detecting the translation direction of parallel sentences, using only an off-the-shelf multilingual NMT system. Experiments on WMT data showed that our approach, without any task-specific supervision, is able to detect the translation direction of NMT-translated sentences with relatively high accuracy. The accuracy increases to 96% if the classifier is provided with at least 10 sentences per document. We also found a robust accuracy for translations by human translators. Finally, we applied our approach to a real-world forensic case and found that it supports the hypothesis that the English book is a forgery.
Future work should explore whether our approach can be improved by mitigating directional bias of the NMT model used. Another open question is to what degree our approach will generalize to document-level translation and to translation with large language models.
@@
Translation strategies:
Our main experiments used academic data from the WMT translation task, where care is taken to ensure that different translation methods are clearly separated: NMT translations did not undergo human post-editing, and human translators were instructed to work from scratch. In practice, parallel documents might have undergone a mixture of translation strategies, which makes it more difficult to predict the accuracy of our approach. Specifically, we found that our approach has less-than-chance accuracy on pre-NMT translations. Applying our approach to web-scale parallel corpus filtering might therefore require additional filtering steps to exclude translations of lower quality.
@@
Acknowledgements
JV and RS acknowledge funding by the Swiss National Science Foundation (project MUTAMUR; no. 213976).
@@ Appendix B Comparison of Models (Document Level)
Appendix C Example for Forensic Dataset
DEEN | ENDE | |
Appendix D Data Statistics
source | target sentences | |||||

### v3

@@ 5.7 Application to Real-World Forensic Case
Finally, we apply our approach to the 86 segment pairs of the plagiarism allegation case. We treat the segments as a single document and classify them with M2M-100 using the document-level approach defined in Section 3.2. We find that, according to the model, it is more probable that the English segments are translations of the German segments than vice versa.
@@
Overall, our analysis supports the hypothesis that German is indeed the language of origin in this real-world dataset (forgery hypothesis; Figure 2). Nevertheless, we recommend that additional evidence of different approaches (automated, manual, and qualitative) should be considered before drawing a final conclusion, given the error rate of 5–21% that we observed in experiments on WMT (ende).
@@ 6 Conclusion
We proposed a novel approach to detecting the translation direction of parallel sentences, using only an off-the-shelf multilingual NMT system. Experiments on WMT data showed that our approach, without any task-specific supervision, is able to detect the translation direction of NMT-produced translations with relatively high accuracy, proving competitive to supervised baselines. Accuracy increases to 96% if the classifier is provided with at least 10 sentences per document. We also found a robust accuracy for translations by human translators. Finally, we applied our approach to a real-world forensic case and found that it supports the hypothesis that the English book is a forgery. Future work should explore whether our approach can be improved by mitigating directional bias of the NMT model used. Another open question is to what degree our approach will generalize to document-level translation and to translation with large language models.
@@
Translation production workflow:
Our main experiments used academic data from the WMT translation task, where care is taken to ensure that different translation methods are clearly separated: NMT translations did not undergo human post-editing, and human translators were instructed to work from scratch. In practice, parallel documents might have undergone a mixture of translation strategies, which makes it more difficult to predict the accuracy of our approach. Specifically, we found that our approach has less-than-chance accuracy on pre-NMT translations. Applying our approach to web-scale parallel corpus filtering Kreutzer et al. (2022); Ranathunga et al. (2024) might therefore require additional filtering steps to exclude translations of lower quality.
@@ Appendix B Comparison of Models (Document Level)
Appendix C Supervised Baseline (WMT) Result Validation Set
HT | NMT | Pre-NMT | ||||||||
Language Pair | Avg. | Avg. | Avg. | Overall Avg. | ||||||
Learning Rate: 1e-05, Epoch 2 | ||||||||||
encs | 83.00 | 89.00 | 86.00 | 87.00 | 89.00 | 88.00 | 89.00 | 89.00 | 89.00 | 87.67 |
ende | 84.00 | 78.00 | 81.00 | 85.00 | 78.00 | 81.50 | 77.00 | 82.00 | 79.50 | 80.67 |
enru | 80.00 | 96.00 | 88.00 | 90.00 | 96.00 | 93.00 | 94.00 | 84.00 | 89.00 | 90.00 |
Macro-Avg. | 82.33 | 87.67 | 85.00 | 87.33 | 87.67 | 87.50 | 86.67 | 85.00 | 85.83 | 86.11 |
Learning Rate: 1e-05, Epoch 4 | ||||||||||
encs | 84.00 | 88.00 | 86.00 | 90.00 | 86.00 | 88.00 | 85.00 | 92.00 | 88.50 | 87.50 |
ende | 92.00 | 73.00 | 82.50 | 93.00 | 74.00 | 83.50 | 71.00 | 93.00 | 82.00 | 82.67 |
enru | 87.00 | 87.00 | 87.00 | 94.00 | 84.00 | 89.00 | 87.00 | 92.00 | 89.50 | 88.50 |
Macro-Avg. | 87.67 | 82.67 | 85.17 | 92.33 | 81.33 | 86.83 | 81.00 | 92.33 | 86.67 | 86.22 |
Learning Rate: 1e-05, Epoch 5 | ||||||||||
encs | 85.00 | 87.00 | 86.00 | 90.00 | 85.00 | 87.50 | 84.00 | 92.00 | 88.00 | 87.17 |
ende | 85.00 | 79.00 | 82.00 | 89.00 | 77.00 | 83.00 | 77.00 | 86.00 | 81.50 | 82.17 |
enru | 75.00 | 97.00 | 86.00 | 86.00 | 95.00 | 90.50 | 94.00 | 79.00 | 86.50 | 87.67 |
Macro-Avg. | 81.67 | 87.67 | 84.67 | 88.33 | 85.67 | 87.00 | 85.00 | 85.67 | 85.33 | 85.67 |
Learning Rate: 2e-05, Epoch 2 | ||||||||||
encs | 81.00 | 88.00 | 84.50 | 82.00 | 87.00 | 84.50 | 87.00 | 84.00 | 85.50 | 84.83 |
ende | 67.00 | 88.00 | 77.50 | 68.00 | 90.00 | 79.00 | 89.00 | 67.00 | 78.00 | 78.17 |
enru | 66.00 | 96.00 | 81.00 | 74.00 | 97.00 | 85.50 | 96.00 | 71.00 | 83.50 | 83.33 |
Macro-Avg. | 71.33 | 90.67 | 81.00 | 74.67 | 91.33 | 83.00 | 90.67 | 74.00 | 82.33 | 82.11 |
Learning Rate: 2e-05, Epoch 4 | ||||||||||
encs | 82.00 | 85.00 | 83.50 | 84.00 | 87.00 | 85.50 | 83.00 | 85.00 | 84.00 | 84.33 |
ende | 86.00 | 71.00 | 78.50 | 89.00 | 74.00 | 81.50 | 72.00 | 85.00 | 78.50 | 79.50 |
enru | 85.00 | 84.00 | 84.50 | 87.00 | 84.00 | 85.50 | 85.00 | 83.00 | 84.00 | 84.67 |
Macro-Avg. | 84.33 | 80.00 | 82.17 | 86.67 | 81.67 | 84.17 | 80.00 | 84.33 | 82.17 | 82.83 |
Learning Rate: 2e-05, Epoch 5 | ||||||||||
encs | 91.00 | 84.00 | 87.50 | 89.00 | 80.00 | 84.50 | 82.00 | 91.00 | 86.50 | 86.17 |
ende | 83.00 | 76.00 | 79.50 | 87.00 | 78.00 | 82.50 | 78.00 | 84.00 | 81.00 | 81.00 |
enru | 83.00 | 85.00 | 84.00 | 85.00 | 88.00 | 86.50 | 86.00 | 85.00 | 85.50 | 85.33 |
Macro-Avg. | 85.67 | 81.67 | 83.67 | 87.00 | 82.00 | 84.50 | 82.00 | 86.67 | 84.33 | 84.17 |
Learning Rate: 3e-05, Epoch 2 | ||||||||||
encs | 81.00 | 91.00 | 86.00 | 81.00 | 89.00 | 85.00 | 89.00 | 82.00 | 85.50 | 85.50 |
ende | 0.00 | 100.00 | 50.00 | 0.00 | 100.00 | 50.00 | 100.00 | 0.00 | 50.00 | 50.00 |
enru | 59.00 | 98.00 | 78.50 | 61.00 | 99.00 | 80.00 | 98.00 | 55.00 | 76.50 | 78.33 |
Macro-Avg. | 46.67 | 96.33 | 71.50 | 47.33 | 96.00 | 71.67 | 95.67 | 45.67 | 70.67 | 66.83 |
Learning Rate: 3e-05, Epoch 4 | ||||||||||
encs | 77.00 | 92.00 | 84.50 | 80.00 | 91.00 | 85.50 | 88.00 | 80.00 | 84.00 | 84.67 |
ende | 100.00 | 0.00 | 50.00 | 100.00 | 0.00 | 50.00 | 0.00 | 100.00 | 50.00 | 50.00 |
enru | 63.00 | 98.00 | 80.50 | 72.00 | 98.00 | 85.00 | 98.00 | 65.00 | 81.50 | 82.33 |
Macro-Avg. | 80.00 | 63.33 | 71.67 | 84.00 | 63.00 | 73.50 | 62.00 | 81.67 | 71.83 | 72.33 |
Learning Rate: 3e-05, Epoch 5 | ||||||||||
encs | 75.00 | 91.00 | 83.00 | 76.00 | 88.00 | 82.00 | 90.00 | 75.00 | 82.50 | 82.50 |
ende | 100.00 | 0.00 | 50.00 | 100.00 | 0.00 | 50.00 | 0.00 | 100.00 | 50.00 | 38.89 |
enru | 69.00 | 98.00 | 83.50 | 75.00 | 96.00 | 85.50 | 95.00 | 71.00 | 83.00 | 84.00 |
Macro-Avg. | 81.33 | 63.00 | 72.17 | 83.67 | 61.33 | 72.50 | 61.67 | 82.00 | 71.83 | 72.17 |
Learning Rate: Dynamic, Epoch 2 | ||||||||||
encs | 83.00 | 88.00 | 85.50 | 80.00 | 85.00 | 82.50 | 88.00 | 82.00 | 85.00 | 84.33 |
ende | 81.00 | 83.00 | 82.00 | 79.00 | 83.00 | 81.00 | 85.00 | 79.00 | 82.00 | 81.67 |
enru | 83.00 | 90.00 | 86.50 | 89.00 | 90.00 | 89.50 | 90.00 | 82.00 | 86.00 | 87.33 |
Macro-Avg. | 82.33 | 87.00 | 84.67 | 82.67 | 86.00 | 84.33 | 87.67 | 81.00 | 84.33 | 84.44 |
Learning Rate: Dynamic, Epoch 4 | ||||||||||
encs | 83.00 | 89.00 | 86.00 | 85.00 | 90.00 | 87.50 | 88.00 | 83.00 | 85.50 | 86.33 |
ende | 88.00 | 75.00 | 81.50 | 93.00 | 78.00 | 85.50 | 78.00 | 83.00 | 80.50 | 82.50 |
enru | 79.00 | 88.00 | 83.50 | 87.00 | 88.00 | 87.50 | 90.00 | 81.00 | 85.50 | 85.50 |
Macro-Avg. | 83.33 | 84.00 | 83.67 | 88.33 | 85.33 | 86.83 | 85.33 | 82.33 | 83.83 | 84.78 |
Learning Rate: Dynamic, Epoch 5 | ||||||||||
encs | 75.00 | 92.00 | 83.50 | 77.00 | 90.00 | 83.50 | 90.00 | 80.00 | 85.00 | 84.00 |
ende | 88.00 | 77.00 | 82.50 | 88.00 | 78.00 | 83.00 | 77.00 | 82.00 | 79.50 | 81.67 |
enru | 79.00 | 91.00 | 85.00 | 87.00 | 88.00 | 87.50 | 89.00 | 79.00 | 84.00 | 85.50 |
Macro-Avg. | 80.67 | 86.67 | 83.67 | 84.00 | 85.33 | 84.67 | 85.33 | 80.33 | 82.83 | 83.72 |
Appendix D Supervised Baseline (WMT) Result Validation Set with Joint Encoding
LP | Input Order | HT | NMT | ||||
Avg. | Avg. | ||||||
ende | source + translation | 79.51 | 65.45 | 72.48 | 83.13 | 69.23 | 76.18 |
translation + source | 78.59 | 60.86 | 69.72 | 83.91 | 65.57 | 74.74 | |
enru | source + translation | 66.13 | 80.73 | 73.43 | 70.31 | 81.88 | 76.1 |
translation + source | 67.64 | 83.24 | 75.44 | 69.57 | 84.05 | 76.81 | |
encs | source + translation | 61.26 | 69.3 | 65.28 | 62.24 | 69.65 | 65.94 |
translation + source | 61.7 | 66.89 | 64.3 | 63.15 | 67.51 | 65.33 |
Model and Data
In addition to the siamese architecture outlined in Section 4.2, which encodes each text segment independently before concatenating their respective embeddings, we also investigate an alternative approach that involves joint encoding of the source and translation texts. Specifically, the two text segments are concatenated with a special delimiter token (“</s>”) inserted between them and then passed together through the encoder Ranasinghe et al. (2020); Rei et al. (2022). The joint encoding system is trained using the same hyperparameters as outlined in Section 4.2, ensuring a consistent comparison.
To mitigate potential input order bias, we reversed the input order for half of the training data. However, this strategy did not entirely eliminate the bias. Consequently, we report test results for both input directions in Table 16 to provide a comprehensive evaluation.
Ultimately, we observed that the siamese architecture demonstrated greater stability with respect to input order, without incurring a significant performance trade-off. Based on these findings, we focus on the siamese architecture as the main baseline presented in this paper.
Appendix E Supervised Baseline (Europarl) and Cross-Domain Application
LP | Approach | HT (Europarl) | HT (WMT) | NMT (WMT) | ||||||
Avg. | Avg. | Avg. | ||||||||
encs | sup (WMT) | — | — | — | 64.70 | 71.36 | 68.03 | 65.43 | 71.98 | 68.71 |
sup (Europarl) | 71.51 | 83.14 | 77.32 | 82.03 | 48.39 | 65.21 | 82.72 | 50.18 | 66.45 | |
unsup | — | — | — | 68.85 | 65.19 | 67.02 | 71.87 | 78.30 | 75.09 | |
ende | sup (WMT) | — | — | — | 87.05 | 55.29 | 71.17 | 89.52 | 61.08 | 75.30 |
sup (Europarl) | 82.12 | 70.80 | 76.46 | 80.97 | 45.36 | 63.17 | 84.15 | 39.70 | 61.93 | |
unsup | — | — | — | 56.38 | 67.44 | 61.91 | 62.69 | 85.27 | 73.98 | |
enru | sup (WMT) | — | — | — | 56.65 | 81.74 | 69.19 | 61.72 | 82.97 | 72.35 |
sup (Europarl) | — | — | — | — | — | — | — | — | — | |
unsup | — | — | — | 71.81 | 54.05 | 62.93 | 76.91 | 71.98 | 74.44 | |
defr | sup (WMT) | — | — | — | — | — | — | — | — | — |
sup (Europarl) | 59.80 | 56.67 | 58.24 | 58.28 | 52.78 | 55.53 | 58.85 | 53.38 | 56.11 | |
unsup | — | — | — | — | — | — | — | — | — |
Language Pair | HT (Europarl) | HT (WMT) | NMT (WMT) |
defr | 0.031 | 0.055 | 0.055 |
csen | 0.116 | 0.336 | 0.325 |
deen | 0.113 | 0.356 | 0.444 |
Macro-Avg. | 0.087 | 0.249 | 0.273 |
Model and Data
In addition to the supervised system trained on WMT16 data, we train a supervised system on a subset of Europarl Koehn (2005); Ustaszewski (2019), a corpus consisting of parallel text from the proceedings of the European Parliament. We train a system for each language pair defr, csen, deen as described in Subsection 4.2. We train on 10,249 samples per direction, use a validation and test set of 1,281 further samples per direction each. Additionally, we apply the systems to the corresponding WMT test data subset as in Subsection 4.2, but with WMT22 defr instead of WMT22/23 enru, to test the systems’ cross-domain capabilities. We experiment in the same manner with several learning rates and epochs as in Appendix C and select the best performing system for each language pair based on the validation accuracy (defr: lr = dynamic, epoch = 5; csen: lr = dynamic, epoch = 5; deen: lr = 1e-05, epoch = 5).
Results
Table 17 shows the results for the unsupervised approach, as well as both the WMT- and Europarl-based supervised systems. While the systems reach higher in-domain accuracies for encs and ende than the WMT-based system, for defr the scores are substantially lower. The results decrease further when the system is tested on out-of-domain data, namely the WMT newstest data. On average, a decrease of 9.37% for WMT human translations and 9.18% NMT-based WMT translations is observable. These results on cross-domain applicability for supervised translation direction detection align with previous work Sominsky and Wintner (2019). A further notable observation, when comparing the in-domain to out-of-domain results, is the increased directional bias, which triples when the systems are applied to out-of-domain data (Table 18).
Appendix F Comparison of Results for Individual WMT Datasets
M2M-100-418M | SMaLL-100 | NLLB-200-1.3B | |||||||
Language Pair | WMT16 | WMT22 | WMT23 | WMT16 | WMT22 | WMT22 | WMT16 | WMT22 | WMT23 |
encs | 66.43 | 73.44 | 71.77 | 64.44 | 71.76 | 67.57 | 50.72 | 67.09 | 62.77 |
ende | 61.88 | 71.65 | - | 63.34 | 71.53 | - | 46.53 | 70.27 | - |
enru | 63.91 | 73.34 | 74.05 | 62.29 | 72.28 | 72.83 | 44.30 | 65.20 | 70.60 |
enuk | - | 78.98 | 74.38 | - | 77.34 | 73.68 | - | 67.53 | 67.85 |
enzh | - | 76.18 | 75.60 | - | 75.97 | 75.52 | - | 51.97 | 58.87 |
csuk | - | 75.39 | 72.09 | - | 75.47 | 76.63 | - | 72.89 | 70.08 |
defr | - | 69.58 | - | - | 71.82 | - | - | 68.48 | - |
Macro-Avg. | 64.08 | 74.08 | 73.58 | 63.36 | 73.74 | 73.25 | 47.19 | 66.20 | 66.03 |
Results by WMT year and model
The WMT16 test data predates the release of all models in our experiments. Additionally, the WMT22 data was released before the NLLB model. This circumstance raises the possibility of data contamination, as the models may have been exposed to test data during training. To address this concern, we compare results across different WMT datasets and add test samples from WMT23, which was released after all three models. If the accuracies for a dataset predating the release of a model are noticeably higher than for a more recent dataset, it would suggest data contamination. However, as shown in 19, this is not observed, indicating that our results are likely unaffected by data contamination.
Appendix G Data Statistics
Source | Reference | |||||
@@ Appendix G Data Statistics
Appendix H Sentence Length
Language Pair | 0-19 | 20-39 | 40-59 | 60-79 | 80-99 | 100-119 | 120-139 | 140-159 | 160-179 | 180-199 |
NMT en-cs | ||||||||||
NMT en-de | ||||||||||
NMT en-ru | ||||||||||
NMT en-uk | ||||||||||
NMT en-zh | ||||||||||
NMT cs-uk | ||||||||||
NMT de-fr | ||||||||||
Macro Avg. |
Appendix I Example for Forensic Dataset
DEEN | ENDE | |